10) What is the magnitude of acceleration of a speck of clay on the edge of a potters wheel turning at 45 rpm (revolution per minute) if the wheels radius is 35 cm? A) 9.78 m/s B) 9.80 m/s C) 10.20 m/s D) 7.78 m/s2 E) 12.40 m/s
Answer #1

10)

w = angular velocity of the wheel = 45 rpm = 45 (2\pi) rad /(60) sec = 4.71 rad/s            Since 1 revolution = 2\pi rad            and 1 min = 60 sec

r = radius of the wheel = 35 cm = 0.35 m

magnitude of acceleration is given as

a = rw2

inserting the values

a = (0.35) (4.71)2

a = 7.78 m/s2